# Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

### v3

@@ 6 Discussion
More Methods That Potentially Improve Accuracy while Reducing COBias. Prior methods tackle the prediction bias problem from different angles. Although COBias is not be explicitly modeled, they are promising to reduce COBias. Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations. In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [36] to estimate a calibration term using test examples.
Figure 8: Comparisons to more methods.
Figure 8 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT and GPT-4 [38] are interesting to evaluate, they do not return logits/softmax probabilities over the entire vocabulary at an output token. For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
@@ 7 Conclusion
This work tackles the imbalance in class prediction accuracy of language models. We introduce COBias and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as an inference-time combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions. For future works, we may extend our studies to more tasks and other modalities. We aim to find simple inference-time plug-in solution for further boosting language model prediction performances.
@@ Appendix B Directly Optimizing the Odd Prediction Rate
We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq. 4), ZCBs (Eq. 14), and ZH (Eq. 17). Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 7 to 13.
@@ Appendix C Derivation of SA Algorithm Complexity
Appendix D Additional Heatmaps for Pairwise Accuracy Differences
Figure 9 and 10 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
@@ Appendix D Additional Heatmaps for Pairwise Accuracy Differences
Appendix E Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 11 and 12.
@@ Appendix E Additional Ablation Analysis
Appendix F Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 14.

### v5

@@ 6 Discussion
More Methods That Potentially Improve Accuracy while Reducing COBias. Prior methods tackle the prediction bias problem from different angles. Although COBias is not be explicitly modeled, they are promising to reduce COBias. Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations. In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [29] to estimate a calibration term using test examples. Figure 6 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT [30] are interesting to debias, ChatGPT returns log probabilities of top 20 most likely tokens, which may not cover the probability for every class that is needed as inputs to DNIP.
Appendix H has more discussions on optimization algorithms for integer programming, how our motivation is different from other classical post-hoc correction methods, using calibration methods to mitigate ICL biases, and using Chain-of-Thought prompting to improve LLM predictions.
@@ 7 Conclusion
We introduce COBias to quantify class accuracy imbalance for language models, and emphasize the need to minimize COBias to push for more reliable predictions while maintaining good overall accuracy. We correct ICL output class probabilities towards this goal as a combinatorial optimization problem, proposing DNIP to optimize class correction weights. Evaluations on three LLMs across seven classification tasks demonstrate that DNIP achieves both lower COBias and higher overall accuracy over original ICL results.
@@ Appendix B Directly Optimizing the Odd Prediction Rate
We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq. 3), ZCBs, and ZH (Eq. 12). Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 8 to 14.
@@ Appendix C Derivation of SA Algorithm Complexity
In addition, the brute force strategy to find optimal correction weights is enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not finish in polynomial time. When $M$ or $N$ is large, enumerated search is computationally intensive. Considering this, we did not adopt the brute force strategy.
Appendix D Dataset Preprocessing
We take the full training set and obtain ICL class probabilities with ground-truth labels for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples, since the datasets are quite large. We further split the training set into 95% optimization set and 5% development set. Test evaluations are performed on each test set against overall accuracy and COBias, except for AGNews, for which we evaluate on 5,000 randomly selected test examples.
Appendix E Performance Improvements vs. Number of Optimization Examples
Figure 8: Relative test COBias reduction; higher is better.
Figure 9: Relative test accuracy increase; higher is better.
Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix I.
Appendix F Additional Heatmaps for Pairwise Accuracy Differences
Figure 10 and 11 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
@@ Appendix F Additional Heatmaps for Pairwise Accuracy Differences
Appendix G Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 12 and 13.
@@ Appendix G Additional Ablation Analysis
Appendix H More Discussions
Appendix I Full Table with Varying Number of Optimization Examples for DNIP
As an addition to Appendix E, the full results for all three models across varying optimization set sizes are shown in Table 15.